OpenStudy (anonymous):

whats the square root of 5 in radical form?

OpenStudy (lalaly):

\[\sqrt{5}\] :S

OpenStudy (anonymous):

?

OpenStudy (anonymous):

just put a root bracket around 5 no?

OpenStudy (lalaly):

yeah its the same, it is in radical form

OpenStudy (anonymous):

oh... im must be confused. i have this problem..

OpenStudy (lalaly):

\[5\sqrt{112}=\sqrt{16 \times 7} =5 \times4\sqrt 7\]

OpenStudy (anonymous):

i have to put it in the radical form

OpenStudy (lalaly):

i forgot a 5 after the first equal sign

OpenStudy (anonymous):

oh wait, haha nevermind.